# Federated Learning with Dynamic Client Arrival and Departure: Convergence and Rapid Adaptation via Initial Model Construction

### v2

@@ 5 Dynamic Initial Model Construction for Fast Adaptation
Pilot Model: As defined in 13, the pilot model is established upon the completion of the pilot preparation stage (the first $P$ sessions). It is calculated as the simple average of the final global model from each of these $P$ sessions:
$\bm{w}_{\text{Pilot}}=\frac{\sum_{s=0}^{P-1}\bm{w}^{(sT+T-1)}}{P}\triangleq\frac{\sum_{s=0}^{P-1}\bm{w}_{s}^{\text{last}}}{P}$ (3)
We introduce the notation $\bm{w}_{s}^{\text{last}}\triangleq\bm{w}^{(sT+T-1)}$ to concisely represent the global model saved at the end of session $s$. These $P$ saved global models are subsequently discarded to free memory.
Gradient Computation and Initial Model Construction: Following the pilot model preparation stage, each subsequent session $s$ starts with a gradient computation process designed to capture the underlying data distribution. This process uses the pilot model as the initial model, $\bm{w}_{s}^{(0)}=\bm{w}_{\text{Pilot}}$, and runs V additional global rounds to obtain a final model. $\bm{w}_{s}^{(V-1)}$ (2 to 5). The computed gradient $G_{s}$ is then defined as the difference $\bm{w}_{s}^{(V-1)}-\bm{w}_{\text{Pilot}}$ (7):
$G_{s}\triangleq\bm{w}_{s}^{(V-1)}-\bm{w}_{s}^{(0)}=\bm{w}_{s}^{(V-1)}-\bm{w}_{\text{Pilot}}.$ (4)
This quantity is particularly useful for evaluating the similarity between different data distributions, and its value is independent of client data amount. For each session following the pilot preparation stage, only one computed gradient is retained. If at least one such gradient has been saved after the pilot stage (6), the initial model for session $s$ is constructed using only the models saved after the pilot stage.
$\bm{w}_{s}^{\text{init}}\triangleq\sum\nolimits_{z=P}^{s-1}\alpha_{s,z}\bm{w}_{z}^{\text{last}},\quad\alpha_{s,z}=\frac{e^{-R\|G_{s}-G_{z}\|_{2}}}{\sum_{z=P}^{s-1}e^{-R\|G_{s}-G_{z}\|_{2}}}.$ (5)
The update rule in (5) defines a weighted average of the saved models $\bm{w}_{z}^{\text{last}}$, where $\alpha_{s,z}$ denotes the corresponding weight. A small difference between the past computed gradient $G_{z}$ and the current gradient $G_{s}$ suggests that the data distribution in the $z$-th session is similar to that of the current session. Since the goal of each session is to learn a model that performs well on the current data distribution, the corresponding saved model $\bm{w}_{z}^{\text{last}}$ serves as a suitable initialization. In such cases, a smaller $\|G_{s}-G_{z}\|_{2}$ yields a larger $\alpha_{s,z}$, thereby assigning more weight to $\bm{w}_{z}^{\text{last}}$ in (5) and making it a dominant component of the constructed initial model.
The constant $R$ in $\alpha_{s,z}$ controls the sensitivity to gradient differences. When $R=0$, the initial model reduces to the simple average of all saved models after the pilot preparation stage, regardless of data similarity. In contrast, as $R\rightarrow\infty$, the initial model converges to the saved model associated with the most similar data distribution.
Remark 1.
Our dynamic initial model construction algorithm provides a simple, effective, and easily integrated approach compatible with any federated learning method. Although providing formal guarantees for warm-start strategies is notoriously challenging [33, 34], especially under client heterogeneity and churn, preliminary theoretical insights (Theorem 1) and extensive empirical validation in Section 6 demonstrate its practical utility. To keep the main text focused, we defer detailed discussions to the appendix: Appendix B covers gradient similarity, sensitivity experiments on $R$ and how well computed gradients capture the underlying data distribution, as well as ablation studies; Appendix C analyzes computation, communication and storage costs.
6 Experiments
FL Algorithm and Baselines: We evaluate the proposed initial model construction method by integrating it with established FL algorithms, including FedAvg [11], FedProx [12], SCAFFOLD [13] and FedACG [14]. Notably, FedACG represents a state-of-the-art approach that has demonstrated superior performance, surpassing methods such as FedDyn [35]. For comparative analysis, we also consider three baseline approaches: Previous, Average, and Continuous (for image dataset). The Previous baseline utilizes the model from the preceding session for continued training, foregoing the explicit construction of a dedicated initial model [36, 37, 11, 38]. The Average baseline initializes models using the unweighted average of prior global models. The Continuous baseline employs the Class-Semantic Relation Distillation Loss [29], which uses the model from the previous session as an anchor for the current client local model. This encourages the new model to preserve relationships between class representations from prior sessions while adapting to incoming client data.
Table 1: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for CIFAR datasets.
Dataset Overlap Dir($\alpha$) FL Alg. First Session Transition Second Session Transition
CIFAR10 0.0 0.3 FedAvg 65.52±2.77 68.96±2.12 69.60±2.51 74.63±1.67 63.85±5.30 61.25±5.54 47.61±1.54 64.37±4.06
FedProx 69.88±3.95 69.12±2.22 71.43±2.77 74.63±2.13 68.07±3.28 61.19±5.87 52.76±2.63 67.66±2.70
SCAFFOLD 80.49±1.08 76.39±2.37 74.21±1.45 80.13±1.31 77.25±1.05 68.60±4.61 72.88±0.83 72.79±1.11
FedACG 80.10±3.69 69.03±2.41 35.46±7.74 51.24±8.33 73.35±4.07 61.65±5.26 32.21±2.56 56.68±6.25
0.7 FedAvg 79.15±1.31 77.33±0.66 76.04±2.09 79.91±0.36 74.63±2.55 70.22±2.81 50.92±1.53 70.63±1.89
FedProx 79.32±1.32 77.35±0.58 77.70±0.48 80.77±0.55 76.29±0.93 70.50±2.54 56.36±0.81 72.05±1.44
SCAFFOLD 84.62±0.36 82.96±0.03 79.59±0.80 83.43±0.38 81.15±0.90 71.46±5.91 75.57±2.02 74.88±1.05
FedACG 86.33±1.22 77.49±0.70 40.10±5.43 62.38±9.57 82.04±2.98 70.40±2.41 36.22±1.92 58.53±9.42
0.2 0.3 FedAvg 61.96±7.12 63.19±3.24 67.95±2.66 69.11±1.35 62.81±5.75 61.23±3.19 57.71±3.46 61.58±1.24
FedProx 67.42±5.09 63.69±3.15 68.76±1.77 68.99±1.37 67.07±3.80 60.94±4.36 60.95±4.89 64.41±2.09
SCAFFOLD 76.45±1.20 70.17±4.42 69.76±2.06 76.31±2.25 77.39±0.94 67.34±3.10 71.34±3.17 72.76±0.35
FedACG 74.96±2.18 63.57±3.40 39.80±2.28 57.44±7.00 76.82±2.60 59.94±4.85 36.45±2.19 59.72±3.47
0.7 FedAvg 74.17±3.36 71.04±0.62 73.42±0.51 74.22±0.82 73.14±4.08 70.57±1.29 64.26±0.87 67.37±0.63
FedProx 75.71±0.30 70.91±0.67 73.83±0.61 73.76±1.28 76.05±1.53 70.43±1.41 66.23±1.13 69.14±1.65
SCAFFOLD 80.21±0.38 76.19±0.68 75.92±0.42 79.58±0.62 80.75±1.08 73.19±3.29 76.24±1.06 76.29±1.36
FedACG 80.68±0.40 70.99±0.54 47.62±8.73 71.47±8.01 81.66±0.58 70.52±1.35 38.88±5.44 71.23±2.63
CIFAR 100 0.0 0.3 FedAvg 47.87±1.26 41.71±1.20 44.60±0.91 47.33±1.09 51.33±0.82 42.88±1.59 45.39±1.93 41.50±3.71
FedProx 47.09±0.54 41.89±0.93 45.80±1.23 47.73±0.18 50.92±0.53 43.20±1.37 48.81±0.84 44.11±1.10
SCAFFOLD 51.68±1.00 44.51±1.14 47.77±1.35 52.35±0.64 54.64±1.57 43.34±0.45 50.74±2.09 49.16±1.10
FedACG 45.83±1.52 41.61±0.32 3.82±2.15 4.01±1.80 48.33±5.25 46.63±1.27 2.00±0.40 10.67±8.36
0.7 FedAvg 52.14±0.29 46.18±0.56 47.83±0.25 51.04±0.59 54.60±0.87 45.78±1.61 48.30±0.75 46.65±1.41
FedProx 50.77±0.70 46.17±1.30 49.77±0.43 51.32±0.64 54.25±0.69 46.56±0.26 52.34±0.72 47.37±0.99
SCAFFOLD 55.00±0.09 52.25±0.79 52.01±0.88 55.66±0.55 57.60±0.27 48.93±1.36 54.52±0.91 51.63±1.06
FedACG 52.97±4.84 46.29±1.05 1.82±1.49 5.91±6.76 55.38±6.27 50.71±0.46 2.59±1.47 22.75±15.49
0.2 0.3 FedAvg 46.95±1.24 42.03±0.28 44.78±0.61 46.11±0.63 50.29±1.53 43.41±0.94 47.26±2.04 42.89±0.67
FedProx 46.35±0.51 42.21±0.16 45.08±0.73 47.24±0.48 50.61±0.16 43.74±1.19 48.53±0.80 44.06±0.20
SCAFFOLD 50.20±1.16 45.94±0.79 48.23±0.49 51.01±0.84 54.04±0.73 45.23±0.73 52.18±0.84 49.08±0.27
FedACG 56.46±0.94 42.83±0.70 20.87±0.28 34.79±5.97 57.58±0.28 44.27±0.33 22.31±2.51 41.11±4.48
0.7 FedAvg 50.56±0.28 46.62±0.23 48.28±0.52 48.64±0.52 54.05±0.70 48.38±0.15 52.18±0.13 45.85±1.05
FedProx 48.90±0.57 46.60±0.41 48.33±1.52 50.24±0.42 53.12±0.82 48.24±0.41 52.41±0.83 47.99±0.95
SCAFFOLD 53.27±0.68 52.21±0.25 50.87±0.91 54.09±0.21 56.76±0.77 51.71±0.42 55.09±0.65 51.86±0.28
FedACG 59.11±0.72 46.66±0.40 23.82±0.84 35.15±9.77 60.70±0.39 48.22±0.06 24.38±1.71 47.24±0.42
Task and Dataset: We perform classification tasks on seven diverse benchmarks, including six image datasets (MNIST [39], Fashion-MNIST [40], SVHN [41], CIFAR-10 [42], CIFAR-100 [42], TinyImageNet [42]) and the AG News text corpus [43]. Results for CIFAR-10, CIFAR-100, and AG News are shown in the main text. Additional experiments appear in Appendix A.
Label Distribution: In our main experiments, we utilize 20 clients, while results with more clients (e.g. 50 or 100) are provided in Appendix A and the supplementary materials. Data samples are distributed among clients by defining the set of labels available within a training session. To introduce variation in the label distribution across the training process, the set of labels available in one training session can partially overlap with the set of labels available in a subsequent session. This overlap level can range from zero (meaning the label sets for consecutive sessions are disjoint) up to a value approaching one (meaning most labels are shared, with only a small fraction unique to each set). For the labels available in a session, they are distributed among the active clients using a Dirichlet distribution parameterized by $\alpha$ [11, 44, 45, 46, 47]. A smaller value of $\alpha$ results in a more heterogeneous distribution of labels across clients, leading to higher non-IIDness between clients. Conversely, a larger value of $\alpha$ yields a more homogeneous distribution of labels across clients.
Table 2: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for AG News text corpus.
Dataset Overlap Dir($\alpha$) FL Alg. First Session Transition Second Session Transition
Proposed Previous Average Proposed Previous Average
AG News 0.0 0.3 FedAvg 68.34±0.34 14.96±0.17 68.94±0.45 73.96±0.84 14.09±0.30 14.30±0.04
FedProx 68.29±0.37 14.95±0.17 68.90±0.46 73.93±0.85 14.08±0.31 14.29±0.03
SCAFFOLD 61.84±0.45 13.13±0.44 60.85±2.03 66.68±1.80 12.96±0.12 12.96±0.34
FedACG 74.71±0.23 16.16±0.21 58.95±1.50 79.91±0.20 15.14±0.10 26.21±1.36
0.7 FedAvg 68.37±0.38 14.95±0.17 68.96±0.48 73.93±0.81 14.11±0.31 14.30±0.03
FedProx 68.32±0.38 14.95±0.16 68.92±0.46 73.89±0.86 14.11±0.31 14.28±0.04
SCAFFOLD 61.90±0.45 13.14±0.43 60.90±2.19 66.73±1.86 12.97±0.14 12.98±0.34
FedACG 74.74±0.21 16.16±0.21 58.96±1.57 79.88±0.25 15.14±0.11 26.18±1.36
0.6 0.3 FedAvg 57.94±0.93 52.12±1.33 51.27±0.93 60.53±0.87 52.54±1.24 52.26±0.91
FedProx 57.86±0.93 52.10±1.33 51.22±0.90 60.45±0.91 52.50±1.24 52.23±0.91
SCAFFOLD 47.88±1.57 47.65±1.57 45.27±1.25 49.51±1.06 49.29±1.49 46.74±1.09
FedACG 67.24±0.91 55.26±0.72 57.44±0.76 70.28±0.88 55.11±1.10 57.69±0.68
0.7 FedAvg 57.88±0.93 52.11±1.27 51.26±0.91 60.50±0.93 52.53±1.25 52.28±0.92
FedProx 57.82±0.95 52.09±1.28 51.22±0.88 60.45±0.94 52.51±1.26 52.25±0.92
SCAFFOLD 48.10±1.49 47.73±1.56 45.34±1.24 49.59±1.07 49.33±1.41 46.83±1.13
FedACG 67.24±0.87 55.21±0.72 57.45±0.76 70.34±0.95 55.10±1.12 57.69±0.73
Results and Takeaways: All values are reported as mean $\pm$ sample standard deviation (%) over three random seeds. Table 1 and Table 2 compare the mean accuracy ($\pm$ std) of our proposed warm start algorithm combined with several federated learning methods against the baseline aggregation schemes in the initial rounds following each session transition on CIFAR10, CIFAR100 and AG News. Across diverse datasets and label distribution scenarios, our approach consistently mitigates the accuracy drop induced by dynamic client arrivals and departures and accelerates recovery whenever the new distribution closely resembles one encountered previously. Figure 2 shows the performance of our proposed algorithm for selected setups across different datasets. It visualizes the fast recovery and robustness client arrivals and departures of our proposed algorithm.
Additional Results: We additionally evaluate our approach with a larger number of clients (e.g., 50 or 100) and on more complex datasets such as TinyImageNet in Appendix A and supplementary materials to further demonstrate its effectiveness. We also examine different label distributions and analyze the impact of the scaling factor $R$ in the appendix and supplementary material.
Figure 2: Performance comparison of the proposed algorithm implemented with selected FL algorithms across selected datasets. The results demonstrate the robustness of our proposed scheme to dynamic data distributions caused by client arrivals and departures.
7 Conclusion
This work addresses the challenges posed by dynamic client participation in FL, where clients may join or leave the system during training. We conduct a theoretical analysis under non-convex loss settings, revealing how factors such as stochastic gradient noise, local training iterations, and non-IID data distributions influence performance. Building on this analysis, we propose a plug-and-play algorithm that constructs an effective initial model for each training round using a weighted averaging strategy guided by gradient similarity. The approach integrates seamlessly with existing FL methods, prioritizing models from clients with data characteristics similar to the active clients, thereby enhancing adaptability and mitigating performance degradation. Experimental results demonstrate significant improvements across diverse datasets and client arrival and departure patterns, underscoring the practical utility of the method in real-world FL scenarios. Future work could refine the initial model construction process to trigger updates only when necessary, reducing computational overhead while maintaining performance and enabling more efficient, adaptive FL systems.
Appendix A Extended Results
A.1 Supplementary Results on Label Distributions from the Main Text
A.1.1 20 clients
Table 3 further illustrates the effectiveness of Algorithm 1 on MNIST, Fashion-MNIST and SVHN dataset for a system of 20 clients.
Table 4 further illustrates the effectiveness of Algorithm 1 on TinyImagenet dataset for a system of 20 clients.
Table 3: Performance comparison of implementing Algorithm 1 with FedProx, FedAvg, SCAFFOLD and FedACG under different label distributions for MNIST, Fashion-MNIST (FMNIST), and SVHN dataset for 20 clients.
Dataset Overlap Dirichlet

### v3

@@ VI-5 Local Model Computation
Ablation Study (Different Pilot Construction Methods): To validate the robustness of our pilot construction strategy, we investigate whether a more sophisticated aggregation method improves the quality of the generated initial models. Specifically, we compare our proposed simple average (Eq. 16) against an Exponentially Weighted Moving Average (EMA) construction. While the simple average assigns uniform importance ($1/P$) to all pilot sessions, the EMA approach prioritizes the most recent global models. Under the EMA scheme, the pilot model is computed as
$\bm{w}_{\text{Pilot}}^{\text{EMA}}=\sum_{s=0}^{P-1}\gamma_{s}\bm{w}_{s}^{\text{last}},$ (29)
where the weighting coefficients $\gamma_{s}$ are determined by a decay factor $\beta\in[0,1)$ as follows:
$\gamma_{s}=\begin{cases}\beta^{P-1},&\text{if }s=0\text{ (oldest session),}\\ (1-\beta)\beta^{P-1-s},&\text{if }0<s\leq P-1.\end{cases}$ (30)
where $\sum_{s=0}^{P-1}\gamma_{s}=1$. For example, with $P=3$ pilot sessions, the aggregated pilot model is computed as: $\bm{w}_{\text{Pilot}}^{\text{EMA}}=\beta^{2}\bm{w}_{0}^{\text{last}}+\beta(1-\beta)\bm{w}_{1}^{\text{last}}+(1-\beta)\bm{w}_{2}^{\text{last}}.$ This pilot formulation theoretically adapts faster to recent trends in the data distribution.
Table IV presents the accuracy comparison across four distinct datasets and FL algorithms. We observe that the performance difference between the two methods is negligible across all scenarios. This consistency implies that the specific weighting of historical models in the pilot phase does not significantly alter the effectiveness of the subsequently computed gradients $G_{s}$. Consequently, we adopt the simple average in Algorithm 1, as it achieves competitive performance without introducing additional hyperparameters.
TABLE V: Ablation Study on the Number of Gradient Computation Rounds ($V$). Comparison of accuracy (mean $\pm$ std) for the Second and Fourth Session Transitions. Using $V=1$ provides the best trade-off between performance and computational efficiency.
Dataset Overlap Dirichlet
$\alpha$
FL Alg. 2nd Session Transition Accuracy (%) 4th Session Transition Accuracy (%)
V=1 V=2 V=3 V=1 V=2 V=3
@@
[…]
Ablation Study (Sensitivity Analysis of $V$): This study evaluates how the number of gradient computation rounds ($V$) affects the balance between performance stability and computational cost. While increasing $V$ theoretically refines gradient estimation, the empirical results in Table V demonstrate that setting $V>1$ yields negligible improvements. Accuracy fluctuations across all datasets remain within standard deviations; for instance, SVHN exhibits a difference of less than $0.03\%$ between $V=1$ and $V=2$. Since computational and communication overheads scale linearly with $V$, we adopt $V=1$ to maximize efficiency without compromising accuracy.
Sensitivity Analysis of Similarity Scaling Factor ($R$): We evaluate the algorithm’s robustness to the scaling factor $R$, which regulates the sensitivity of the weighting coefficients $\mu_{s,z}$ to gradient differences. Table VI summarizes performance across $R$ values tailored to the gradient norms of each dataset (e.g., higher magnitudes for AG News). The results indicate that the algorithm is highly stable across variations in $R$. Accuracy fluctuations are statistically negligible; for instance, SVHN shows a variation of less than $0.1\%$, and the performance margins for AG News and CIFAR-100 largely overlap within standard deviations. This suggests that while $R$ must be set to an appropriate order of magnitude, precise fine-tuning is not necessary to achieve good performance.
TABLE VI: Sensitivity Analysis of the Similarity Scaling Factor ($R$). Comparison of accuracy (mean $\pm$ std) for the Second and Fourth Session Transitions. The algorithm demonstrates robustness to variations in $R$ within reasonable ranges.
Dataset (Values of $R$) Overlap Dirichlet